Hi there! can you please help me answer these questions with steps so I understand how you answered them. This is math statistics, thank you in advance!!
Below are the annual numbers of deaths from tornados in the United States from 1998 to 2016. 23 17 40 40 245 54 35 38 67 81 125 21 65 78 82 95 54 190 69 a) Using your graphing calculator, enter the data into L1. Then go to STAT- CALC- 1-Var Stats and report the 5-number summary. (2.5 points) b) Calculate the fences for a modified boxplot. Be sure to write the formulas and show you work (3 points) Are there any outliers? If so, what are they? (1 point) Construct a modified boxplot and be sure to label each value from the 5-number summary, the fences and any outliers. (5 points) What are the best measures for the center and spread of this distribution? c) d) e) (2 points)
media%2F795%2F795b3fec-5602-46a9-9cbf-47
Answer #1

a)Population size: 19 Median: 65 Minimum: 17 Maximum: 245 First quartile: 38 Third quartile: 82 Interquartile Range: 44 Outliers: 245 190 240 220 200 180 160 140 120 100 80 60 40 20

b) Lower fence is Q_1-1.5IQR =38-1.5*44=-28

Upper fence: Q_3+1.5IQR =82+1.5*44=148

c) Yes there are outlier 245 and 190

d)